# Exploratory Randomization for Discrete-Time Linear Exponential Quadratic Gaussian (LEQG) Problem

### v1

@@ 3.3 Solving the Risk-Neutral Penalized Game Problem
$=$ $\frac{1}{2}\left(a+Ax_{t}+B\bar{u}_{t}\right)^{\prime}P_{t+1}\left(a+Ax_{t}+B\bar{u}_{t}\right)+\frac{1}{2}\text{tr}\left(B^{\prime}\Xi_{t}BP_{t+1}\right)+\frac{1}{2}\bar{\eta}_{t}^{\prime}B^{\prime}P_{t+1}B\bar{\eta}_{t}+\frac{1}{2}\text{tr}\left(\Lambda_{t}P_{t+1}\right)$
$+\frac{1}{2}\bar{\gamma}_{t}^{\prime}P_{t+1}\bar{\gamma}_{t}$
$+\left(a+Ax_{t}+B\bar{u}_{t}\right)^{\prime}P_{t+1}B\bar{\eta}_{t}+\left(a+Ax_{t}+B\bar{u}_{t}\right)^{\prime}P_{t+1}\bar{\gamma}_{t}+\bar{\eta}_{t}^{\prime}B^{\prime}P_{t+1}\bar{\gamma}_{t}$
$+\left(a+Ax_{t}+B\bar{u}_{t}\right)^{\prime}p_{t+1}+\bar{\eta}_{t}^{\prime}B^{\prime}p_{t+1}+\bar{\gamma}_{t}^{\prime}p_{t+1}+r_{t+1}$ (3.18)
@@
$\mathfrak{A}^{(1)}_{t+1}$ $=B^{\prime}P_{t+1}A+\theta Q,$ $\mathfrak{A}^{(2)}_{t+1}$ $=P_{t+1}^{\prime}A$
@@
$\mathfrak{B}^{(1)}_{t+1}$ $=B^{\prime}P_{t+1}B+2\theta N_{t}$ $\mathfrak{B}^{(2)}_{t+1}$ $=-\Lambda^{-1}_{t}+P_{t+1}$ (3.19)
$\mathfrak{B}^{(3)}_{t+1}$ $=-\Xi_{t}^{-1}+\mathfrak{B}^{(1)}_{t+1}=-\Xi_{t}^{-1}+B^{\prime}P_{t+1}B+2\theta N_{t}$
@@
$\mathfrak{G}_{t+1}:=$ $\mathfrak{C}_{t+1}^{\prime}\left(\mathfrak{B}^{(1)}_{t+1}\right)^{-1}\mathfrak{C}_{t+1}-\mathfrak{B}^{(2)}_{t+1}$ (3.20)
@@
$F(\bar{u},\bar{\gamma},\bar{\eta}):=$ $\frac{1}{2}\bar{u}^{\prime}\mathfrak{B}^{(1)}_{t+1}\bar{u}+\bar{u}^{\prime}\left(\mathfrak{A}^{(1)}_{t+1}x_{t}+\mathfrak{a}^{(1)}_{t+1}\right)+\frac{1}{2}\bar{\gamma}^{\prime}\mathfrak{B}^{(2)}_{t+1}\bar{\gamma}+\bar{\gamma}^{\prime}\left(\mathfrak{A}^{(2)}_{t+1}x_{t}+\mathfrak{a}^{(2)}_{t+1}\right)$
$+\frac{1}{2}\bar{\eta}^{\prime}\mathfrak{B}^{(3)}_{t+1}\bar{\eta}+\bar{\eta}^{\prime}\left(\mathfrak{A}^{(1)}_{t+1}x_{t}+\mathfrak{a}^{(1)}_{t+1}\right)+\bar{u}^{\prime}\mathfrak{C}_{t+1}\bar{\gamma}+\bar{u}^{\prime}\mathfrak{B}^{(1)}_{t+1}\bar{\eta}+\bar{\eta}^{\prime}\mathfrak{C}_{t+1}\bar{\gamma},$ (3.21)
@@
$V_{t}(x_{t})=$ $\inf_{\bar{u}}\sup_{\bar{\gamma},\bar{\eta}}\left\{F(\bar{u},\bar{\gamma},\bar{\eta})\right\}+x_{t}^{\prime}\left(\theta M+\frac{1}{2}A^{\prime}P_{t+1}A\right)x_{t}+x_{t}^{\prime}\left(A^{\prime}P_{t+1}a+\theta m+A^{\prime}p_{t+1}\right)$
$+\frac{1}{2}\text{tr}\left(B^{\prime}\Xi_{t}BP_{t+1}\right)+\frac{1}{2}\text{tr}\left(\Lambda_{t}P_{t+1}\right)+r_{t+1}+\frac{1}{2}a^{\prime}P_{t+1}a+a^{\prime}p_{t+1}+\theta\text{tr}\left(\Xi N_{t}\right).$ (3.22)
@@
$u_{t}^{*}$ $=\left(\mathfrak{B}^{(1)^{\prime}}_{t+1}\mathfrak{B}^{(2)}_{t+1}-\mathfrak{C}^{\prime}_{t+1}\mathfrak{C}_{t+1}\right)^{-1}\left[\left(\mathfrak{A}^{(2)^{\prime}}_{t+1}\mathfrak{C}_{t+1}-\mathfrak{A}^{(1)^{\prime}}_{t+1}\mathfrak{B}^{(2)}_{t+1}\right)\,x_{t}+\left(\mathfrak{C}^{\prime}_{t+1}\mathfrak{a}^{(2)}_{t+1}-\mathfrak{B}^{(2)^{\prime}}_{t+1}\mathfrak{a}^{(1)}_{t+1}\right)\right]$
@@
$\hskip 85.35826pt\left.+P^{\prime}_{t+1}BP_{t+1}a+P^{\prime}_{t+1}Bp_{t+1}-\left(P_{t+1}-\Lambda^{-1}\right)\left(B^{\prime}P_{t+1}a+B^{\prime}p_{t+1}+\theta\,n\right)\right\}$
@@
=$\quad=\Bigl{\{}\left[\left(P_{t+1}-\Lambda^{-1}\right)\left(B^{\prime}P_{t+1}B+2\,\theta\,N\right)-B^{\prime}P_{t+1}P_{t+1}B\right]^{-1}$
@@
$\eta_{t}^{*}$ $=-\left(\mathfrak{B}^{(3)}_{t+1}\right)^{-1}\,\left[\mathfrak{A}^{(1^{\prime})}_{t+1}x_{t}+\mathfrak{a}^{(1)}_{t+1}+(\mathfrak{B}^{(1^{\prime})}_{t+1})u_{t}^{*}+\mathfrak{C}_{t+1}\gamma_{t}^{*}\right]$
@@
$V_{t}(x_{t})=$ $\inf_{\bar{u}}\sup_{\bar{\gamma},\bar{\eta}}\left\{\theta x_{t}^{\prime}Mx_{t}+\theta\text{tr}\left(\Xi N_{t}\right)+\theta\bar{u}^{\prime}N_{t}\bar{u}+2\theta\bar{u}^{\prime}N_{t}\bar{\eta}+\theta\bar{\eta}^{\prime}N_{t}\bar{\eta}+\theta\bar{u}^{\prime}Qx_{t}+\theta\bar{\eta}^{\prime}Qx_{t}+\theta\bar{u}^{\prime}n+\theta\bar{\eta}^{\prime}n+\theta x_{t}^{\prime}m\right.$
$\left.-\frac{1}{2}\bar{\gamma}^{\prime}\Lambda^{-1}_{t}\bar{\gamma}-\frac{1}{2}\bar{\eta}^{\prime}\Xi_{t}^{-1}\bar{\eta}+\frac{1}{2}\left(a+Ax_{t}+B\bar{u}\right)^{\prime}P_{t+1}\left(a+Ax_{t}+B\bar{u}\right)+\frac{1}{2}\bar{\gamma}^{\prime}P_{t+1}\bar{\gamma}+\frac{1}{2}\bar{\eta}^{\prime}B^{\prime}P_{t+1}B\bar{\eta}\right.$
$\left.+\left(a+Ax_{t}+B\bar{u}\right)^{\prime}p_{t+1}+\bar{\gamma}^{\prime}p_{t+1}+\bar{\eta}^{\prime}B^{\prime}p_{t+1}+\left(a+Ax_{t}+B\bar{u}\right)^{\prime}P_{t+1}\bar{\gamma}\right.$
$\left.+\left(a+Ax_{t}+B\bar{u}\right)^{\prime}P_{t+1}B\bar{\eta}+\bar{\eta}_{t}^{\prime}B^{\prime}P_{t+1}\bar{\gamma}_{t}+\frac{1}{2}\text{tr}\left(B^{\prime}\Xi_{t}BP_{t+1}\right)+\frac{1}{2}\text{tr}\left(\Lambda_{t}P_{t+1}\right)+r_{t+1}\right\}$
$=$ $\inf_{\bar{u}}\sup_{\bar{\gamma},\bar{\eta}}\left\{\theta x_{t}^{\prime}Mx_{t}+\theta\text{tr}\left(\Xi N_{t}\right)+\theta\bar{u}^{\prime}N_{t}\bar{u}+2\theta\bar{u}^{\prime}N_{t}\bar{\eta}+\theta\bar{u}^{\prime}Qx_{t}+\theta\bar{\eta}^{\prime}Qx_{t}+\theta\bar{u}^{\prime}n+\theta\bar{\eta}^{\prime}n\right.$
@@
$\left.+\frac{1}{2}\text{tr}\left(B^{\prime}\Xi_{t}BP_{t+1}\right)+\frac{1}{2}\text{tr}\left(\Lambda_{t}P_{t+1}\right)+r_{t+1}+\frac{1}{2}a^{\prime}P_{t+1}a+a^{\prime}p_{t+1}+\theta\text{tr}\left(\Xi N_{t}\right)\right\}$ (3.24)
@@
$\mathcal{H}(t,x_{t},\bar{u}_{t},\bar{\gamma}_{t},\bar{\eta}_{t}):=$ $\theta\left[x_{t}^{\prime}Mx_{t}+\text{tr}\left(\Xi N_{t}\right)+\left(\bar{u}+\bar{\eta}\right)^{\prime}N_{t}\left(\bar{u}+\bar{\eta}\right)+\left(\bar{u}+\bar{\eta}\right)^{\prime}Qx_{t}+x_{t}^{\prime}m+\left(\bar{u}+\bar{\eta}\right)^{\prime}n\right]$
$-\frac{1}{2}\left(\bar{\gamma}^{\prime}\Lambda^{-1}_{t}\bar{\gamma}+\bar{\eta}^{\prime}\Xi_{t}^{-1}\bar{\eta}\right)+\mathbf{E}_{t,x_{t}}^{\bar{\gamma},\bar{\eta}}\left[V_{t+1}(x_{t+1})\right]$ (3.25)
@@
$\theta\left[x_{t}^{\prime}Mx_{t}+\text{tr}\left(\Xi N_{t}\right)+\left(u^{*}_{t}+\bar{\eta}_{t}\right)^{\prime}N_{t}\left(u^{*}_{t}+\bar{\eta}_{t}\right)+\left(u^{*}_{t}+\bar{\eta}_{t}\right)^{\prime}Qx_{t}+x_{t}^{\prime}m+\left(u^{*}_{t}+\bar{\eta}_{t}\right)^{\prime}n\right]$
$-\frac{1}{2}\left(\bar{\gamma}_{t}^{\prime}\Lambda^{-1}\bar{\gamma}_{t}+\bar{\eta}^{\prime}\Xi_{t}^{-1}\bar{\eta}_{t}\right)+\mathbf{E}_{t,x_{t}}^{\bar{\gamma}_{t},\bar{\eta}_{t}}\left[V_{t+1}(x_{t+1})\right]$
$\leq$ $\theta\left[x_{t}^{\prime}Mx_{t}+\text{tr}\left(\Xi N_{t}\right)+\left(u^{*}_{t}+\eta^{*}_{t}\right)^{\prime}N_{t}\left(u^{*}_{t}+\eta^{*}_{t}\right)+\left(u^{*}_{t}+\eta^{*}_{t}\right)^{\prime}Qx_{t}+x_{t}^{\prime}m+\left(u^{*}_{t}+\eta^{*}_{t}\right)^{\prime}n\right]$
@@
$\leq$ $\theta\left[x_{t}^{\prime}Mx_{t}+\text{tr}\left(\Xi N_{t}\right)+\left(\bar{u}_{t}+\eta^{*}_{t}\right)^{\prime}N_{t}\left(\bar{u}_{t}+\eta^{*}_{t}\right)+\left(\bar{u}_{t}+\eta^{*}_{t}\right)^{\prime}Qx_{t}+x_{t}^{\prime}m+\left(\bar{u}_{t}+\eta^{*}_{t}\right)^{\prime}n\right]$
$-\frac{1}{2}\left((\gamma^{*}_{t})^{\prime}\Lambda^{-1}\gamma^{*}_{t}+(\eta^{*})^{\prime}\Xi_{t}^{-1}\eta^{*}_{t}\right)+\mathbf{E}_{t,x_{t}}^{\gamma^{*}_{t},\eta^{*}_{t}}\left[V_{t+1}(x_{t+1})\right].$ (3.28)
@@
The next step is to identify the unique stationary point $(u^{*},\gamma^{*},\eta^{*})$ of the quadratic function $F(\bar{u},\bar{\gamma},\bar{\eta})$. By the first order condition, $u^{*}$ satisfies
$\frac{\partial F}{\partial\bar{u}}(\bar{u},\bar{\gamma},\bar{\eta})\Big{|}_{\bar{u}=u^{*}}=\mathfrak{B}^{(1)}_{t+1}u^{*}+\mathfrak{A}^{(1)}_{t+1}x_{t}+\mathfrak{a}^{(1)}_{t+1}+\mathfrak{C}_{t+1}\bar{\gamma}+\mathfrak{B}^{(1)}_{t+1}\bar{\eta}=0$
@@
$\frac{\partial F}{\partial\bar{\eta}}(\bar{u},\bar{\gamma},\bar{\eta})\Big{|}_{\bar{\eta}=\eta^{*}}=\mathfrak{B}^{(3)}_{t+1}\eta^{*}+\mathfrak{A}^{(1)}_{t+1}x_{t}+\mathfrak{a}^{(1)}_{t+1}+(\mathfrak{B}^{(1)}_{t+1})^{\prime}\bar{u}+\mathfrak{C}_{t+1}\bar{\gamma}=0$
@@
$-\left(\mathfrak{B}^{(1)}_{t+1}\right)^{-1}\left[\mathfrak{A}^{(1)}_{t+1}x_{t}+\mathfrak{a}^{(1)}_{t+1}\right]$
@@
$-\left(\mathfrak{B}^{(1)}_{t+1}\right)^{-1}\mathfrak{C}_{t+1}\gamma^{*}$
@@
$\mathfrak{Q}_{t+1}=$ $-\left(\mathfrak{A}^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{A}^{(1)}_{t+1}-2\left(\mathfrak{A}^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{C}_{t+1}(\mathfrak{G}_{t+1})^{-1}\mathfrak{A}^{(2)}_{t+1}$
$+\left(\mathfrak{A}^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{C}_{t+1}(\mathfrak{G}_{t+1})^{-1}\mathfrak{C}_{t+1}^{\prime}\left(\mathfrak{B}^{(1)}_{t+1}\right)^{-1}\mathfrak{A}^{(1)}_{t+1}+\left(\mathfrak{A}^{(2)}_{t+1}\right)^{\prime}(\mathfrak{G}_{t+1})^{-1}\mathfrak{A}^{(2)}_{t+1}$ (3.37)
$\mathfrak{q}_{t+1}=$ $-\left(\mathfrak{A}^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{a}^{(1)}_{t+1}-\left(\mathfrak{A}^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{C}_{t+1}(\mathfrak{G}_{t+1})^{-1}\mathfrak{a}^{(2)}_{t+1}$

### v2

@@ 3.3 Solving the Risk-Neutral Penalized Game Problem
$\mathbf{E}_{t,x_{t}}^{\bar{\gamma},\bar{\eta}}\left[V_{t+1}(x_{t+1})\right]=$ $\frac{1}{2}\left(a+Ax_{t}+B\bar{u}_{t}\right)^{\prime}P_{t+1}\left(a+Ax_{t}+B\bar{u}_{t}\right)$ (3.18)
$\frac{1}{2}\text{tr}\left(B^{\prime}\Xi_{t}BP_{t+1}\right)+\frac{1}{2}\bar{\eta}_{t}^{\prime}B^{\prime}P_{t+1}B\bar{\eta}_{t}+\frac{1}{2}\text{tr}\left(\Lambda_{t}P_{t+1}\right)$
$+\frac{1}{2}\bar{\gamma}_{t}^{\prime}P_{t+1}\bar{\gamma}_{t}+\left(a+Ax_{t}+B\bar{u}_{t}\right)^{\prime}P_{t+1}B\bar{\eta}_{t}$
$+\left(a+Ax_{t}+B\bar{u}_{t}\right)^{\prime}P_{t+1}\bar{\gamma}_{t}+\bar{\eta}_{t}^{\prime}B^{\prime}P_{t+1}\bar{\gamma}_{t}$
$+\left(a+Ax_{t}+B\bar{u}_{t}\right)^{\prime}p_{t+1}+\bar{\eta}_{t}^{\prime}B^{\prime}p_{t+1}+\bar{\gamma}_{t}^{\prime}p_{t+1}+r_{t+1}$
@@
$\mathfrak{A}^{(1)}_{t+1}$ $=B^{\prime}P_{t+1}A+\theta Q,$ $\mathfrak{A}^{(2)}_{t+1}$ $=P_{t+1}^{\prime}A$ (3.19)
@@
$\mathfrak{B}^{(1)}_{t+1}$ $=B^{\prime}P_{t+1}B+2\theta N_{t}$ $\mathfrak{B}^{(2)}_{t+1}$ $=-\Lambda^{-1}_{t}+P_{t+1}$
$\mathfrak{B}^{(3)}_{t+1}$ $=-\Xi_{t}^{-1}+\mathfrak{B}^{(1)}_{t+1}=-\Xi_{t}^{-1}+B^{\prime}P_{t+1}B+2\theta N_{t}$ (3.20)
@@
$\mathfrak{G}_{t+1}$ $=\mathfrak{C}_{t+1}^{\prime}\left(\mathfrak{B}^{(1)}_{t+1}\right)^{-1}\mathfrak{C}_{t+1}-\mathfrak{B}^{(2)}_{t+1}$
@@
$F(\bar{u},\bar{\gamma},\bar{\eta}):=$ $\frac{1}{2}\bar{u}^{\prime}\mathfrak{B}^{(1)}_{t+1}\bar{u}+\bar{u}^{\prime}\left(\mathfrak{A}^{(1)}_{t+1}x_{t}+\mathfrak{a}^{(1)}_{t+1}\right)+\frac{1}{2}\bar{\gamma}^{\prime}\mathfrak{B}^{(2)}_{t+1}\bar{\gamma}$ (3.21)
$+\bar{\gamma}^{\prime}\left(\mathfrak{A}^{(2)}_{t+1}x_{t}+\mathfrak{a}^{(2)}_{t+1}\right)+\frac{1}{2}\bar{\eta}^{\prime}\mathfrak{B}^{(3)}_{t+1}\bar{\eta}+\bar{\eta}^{\prime}\left(\mathfrak{A}^{(1)}_{t+1}x_{t}+\mathfrak{a}^{(1)}_{t+1}\right)$
$+\bar{u}^{\prime}\mathfrak{C}_{t+1}\bar{\gamma}+\bar{u}^{\prime}\mathfrak{B}^{(1)}_{t+1}\bar{\eta}+\bar{\eta}^{\prime}\mathfrak{C}_{t+1}\bar{\gamma},$
@@
$V_{t}(x_{t})=$ $\inf_{\bar{u}}\sup_{\bar{\gamma},\bar{\eta}}\left\{F(\bar{u},\bar{\gamma},\bar{\eta})\right\}+x_{t}^{\prime}\left(\theta M+\frac{1}{2}A^{\prime}P_{t+1}A\right)x_{t}$ (3.22)
$+x_{t}^{\prime}\left(A^{\prime}P_{t+1}a+\theta m+A^{\prime}p_{t+1}\right)+\frac{1}{2}\text{tr}\left(B^{\prime}\Xi_{t}BP_{t+1}\right)+\frac{1}{2}\text{tr}\left(\Lambda_{t}P_{t+1}\right)+r_{t+1}$
$+\frac{1}{2}a^{\prime}P_{t+1}a+a^{\prime}p_{t+1}+\theta\text{tr}\left(\Xi_{t}N_{t}\right).$
@@
$u_{t}^{*}$ $=\left(\mathfrak{B}^{(1)^{\prime}}_{t+1}\mathfrak{B}^{(2)}_{t+1}-\mathfrak{C}^{\prime}_{t+1}\mathfrak{C}_{t+1}\right)^{-1}\left[\left(\mathfrak{A}^{(2)^{\prime}}_{t+1}\mathfrak{C}_{t+1}-\mathfrak{A}^{(1)^{\prime}}_{t+1}\mathfrak{B}^{(2)}_{t+1}\right)\,x_{t}+\left(\mathfrak{C}^{\prime}_{t+1}\mathfrak{a}^{(2)}_{t+1}-\mathfrak{B}^{(2)^{\prime}}_{t+1}\mathfrak{a}^{(1)}_{t+1}\right)\right$] (3.23)
@@
$\left.+P^{\prime}_{t+1}BP_{t+1}a+P^{\prime}_{t+1}Bp_{t+1}-\left(P_{t+1}-\Lambda^{-1}\right)\left(B^{\prime}P_{t+1}a+B^{\prime}p_{t+1}+\theta\,n\right)\right\}$
@@
$\quad=\Bigl\{\left[\left(P_{t+1}-\Lambda^{-1}\right)\left(B^{\prime}P_{t+1}B+2\,\theta\,N\right)-B^{\prime}P_{t+1}P_{t+1}B\right]^{-1}$
@@
$\eta_{t}^{*}$ $=-\left(\mathfrak{B}^{(3)}_{t+1}\right)^{-1}\,\left[\mathfrak{A}^{(1)^{\prime}}_{t+1}x_{t}+\mathfrak{a}^{(1)}_{t+1}+(\mathfrak{B}^{(1)^{\prime}}_{t+1})u_{t}^{*}+\mathfrak{C}_{t+1}\gamma_{t}^{*}\right]$
@@
$V_{t}(x_{t})$
$=$ $\inf_{\bar{u}}\sup_{\bar{\gamma},\bar{\eta}}\left\{\theta x_{t}^{\prime}Mx_{t}+\theta\text{tr}\left(\Xi_{t}N_{t}\right)+\theta\bar{u}^{\prime}N_{t}\bar{u}+2\theta\bar{u}^{\prime}N_{t}\bar{\eta}+\theta\bar{\eta}^{\prime}N_{t}\bar{\eta}+\theta\bar{u}^{\prime}Qx_{t}+\theta\bar{\eta}^{\prime}Qx_{t}+\theta\bar{u}^{\prime}n\right$.
$\left.+\theta\bar{\eta}^{\prime}n+\theta x_{t}^{\prime}m-\frac{1}{2}\bar{\gamma}^{\prime}\Lambda^{-1}_{t}\bar{\gamma}-\frac{1}{2}\bar{\eta}^{\prime}\Xi_{t}^{-1}\bar{\eta}+\frac{1}{2}\left(a+Ax_{t}+B\bar{u}\right)^{\prime}P_{t+1}\left(a+Ax_{t}+B\bar{u}\right)\right$.
$\left.+\frac{1}{2}\bar{\gamma}^{\prime}P_{t+1}\bar{\gamma}+\frac{1}{2}\bar{\eta}^{\prime}B^{\prime}P_{t+1}B\bar{\eta}+\left(a+Ax_{t}+B\bar{u}\right)^{\prime}p_{t+1}+\bar{\gamma}^{\prime}p_{t+1}+\bar{\eta}^{\prime}B^{\prime}p_{t+1}\right$.
$\left.+\left(a+Ax_{t}+B\bar{u}\right)^{\prime}P_{t+1}\bar{\gamma}+\left(a+Ax_{t}+B\bar{u}\right)^{\prime}P_{t+1}B\bar{\eta}+\bar{\eta}_{t}^{\prime}B^{\prime}P_{t+1}\bar{\gamma}_{t}\right$.
$\left.+\frac{1}{2}\text{tr}\left(B^{\prime}\Xi_{t}BP_{t+1}\right)+\frac{1}{2}\text{tr}\left(\Lambda_{t}P_{t+1}\right)+r_{t+1}\right\}$
$=$ $\inf_{\bar{u}}\sup_{\bar{\gamma},\bar{\eta}}\left\{\theta x_{t}^{\prime}Mx_{t}+\theta\text{tr}\left(\Xi_{t}N_{t}\right)+\theta\bar{u}^{\prime}N_{t}\bar{u}+2\theta\bar{u}^{\prime}N_{t}\bar{\eta}+\theta\bar{u}^{\prime}Qx_{t}+\theta\bar{\eta}^{\prime}Qx_{t}+\theta\bar{u}^{\prime}n+\theta\bar{\eta}^{\prime}n\right$.
@@
$\left.+\frac{1}{2}\text{tr}\left(B^{\prime}\Xi_{t}BP_{t+1}\right)+\frac{1}{2}\text{tr}\left(\Lambda_{t}P_{t+1}\right)+r_{t+1}+\frac{1}{2}a^{\prime}P_{t+1}a+a^{\prime}p_{t+1}+\theta\text{tr}\left(\Xi_{t}N_{t}\right)\right\}$ (3.24)
@@
$\mathcal{H}(t,x_{t},\bar{u}_{t},\bar{\gamma}_{t},\bar{\eta}_{t})$ (3.25)
$:=$ $\theta\left[x_{t}^{\prime}Mx_{t}+\text{tr}\left(\Xi N_{t}\right)+\left(\bar{u}+\bar{\eta}\right)^{\prime}N_{t}\left(\bar{u}+\bar{\eta}\right)+\left(\bar{u}+\bar{\eta}\right)^{\prime}Qx_{t}\right$.
$\left.+x_{t}^{\prime}m+\left(\bar{u}+\bar{\eta}\right)^{\prime}n\right]-\frac{1}{2}\left(\bar{\gamma}^{\prime}\Lambda^{-1}_{t}\bar{\gamma}+\bar{\eta}^{\prime}\Xi_{t}^{-1}\bar{\eta}\right)+\mathbf{E}_{t,x_{t}}^{\bar{\gamma},\bar{\eta}}\left[V_{t+1}(x_{t+1})\right]$
@@
$\theta\left[x_{t}^{\prime}Mx_{t}+\text{tr}\left(\Xi_{t}N_{t}\right)+\left(u^{*}_{t}+\bar{\eta}_{t}\right)^{\prime}N_{t}\left(u^{*}_{t}+\bar{\eta}_{t}\right)+\left(u^{*}_{t}+\bar{\eta}_{t}\right)^{\prime}Qx_{t}+x_{t}^{\prime}m+\left(u^{*}_{t}+\bar{\eta}_{t}\right)^{\prime}n\right]$ (3.28)
$-\frac{1}{2}\left(\bar{\gamma}_{t}^{\prime}\Lambda^{-1}\bar{\gamma}_{t}+\bar{\eta}_{t}^{\prime}\Xi_{t}^{-1}\bar{\eta}_{t}\right)+\mathbf{E}_{t,x_{t}}^{\bar{\gamma}_{t},\bar{\eta}_{t}}\left[V_{t+1}(x_{t+1})\right]$
$\leq$ $\theta\left[x_{t}^{\prime}Mx_{t}+\text{tr}\left(\Xi_{t}N_{t}\right)+\left(u^{*}_{t}+\eta^{*}_{t}\right)^{\prime}N_{t}\left(u^{*}_{t}+\eta^{*}_{t}\right)+\left(u^{*}_{t}+\eta^{*}_{t}\right)^{\prime}Qx_{t}+x_{t}^{\prime}m+\left(u^{*}_{t}+\eta^{*}_{t}\right)^{\prime}n\right]$
@@
$\leq$ $\theta\left[x_{t}^{\prime}Mx_{t}+\text{tr}\left(\Xi_{t}N_{t}\right)+\left(\bar{u}_{t}+\eta^{*}_{t}\right)^{\prime}N_{t}\left(\bar{u}_{t}+\eta^{*}_{t}\right)+\left(\bar{u}_{t}+\eta^{*}_{t}\right)^{\prime}Qx_{t}+x_{t}^{\prime}m+\left(\bar{u}_{t}+\eta^{*}_{t}\right)^{\prime}n\right]$
$-\frac{1}{2}\left((\gamma^{*}_{t})^{\prime}\Lambda^{-1}\gamma^{*}_{t}+(\eta^{*})^{\prime}\Xi_{t}^{-1}\eta^{*}_{t}\right)+\mathbf{E}_{t,x_{t}}^{\gamma^{*}_{t},\eta^{*}_{t}}\left[V_{t+1}(x_{t+1})\right].$
@@
The next step is to identify the unique stationary point $(u^{*},\gamma^{*},\eta^{*})$ of the quadratic function $F(\bar{u}_{t},\bar{\gamma},\bar{\eta})$. By the first order condition, $u^{*}$ satisfies
$\frac{\partial F}{\partial\bar{u}}(\bar{u},\bar{\gamma},\bar{\eta})\Big|_{\bar{u}=u^{*}}=0$
$\Leftrightarrow$ $=\mathfrak{B}^{(1)}_{t+1}u^{*}+\mathfrak{A}^{(1)}_{t+1}x_{t}+\mathfrak{a}^{(1)}_{t+1}+\mathfrak{C}_{t+1}\bar{\gamma}+\mathfrak{B}^{(1)}_{t+1}\bar{\eta}=0$
@@
$\frac{\partial F}{\partial\bar{\eta}}(\bar{u},\bar{\gamma},\bar{\eta})\Big|_{\bar{\eta}=\eta^{*}}=0$
$\Leftrightarrow$ $\mathfrak{B}^{(3)}_{t+1}\eta^{*}+\mathfrak{A}^{(1)}_{t+1}x_{t}+\mathfrak{a}^{(1)}_{t+1}+(\mathfrak{B}^{(1)}_{t+1})^{\prime}\bar{u}+\mathfrak{C}_{t+1}\bar{\gamma}=0$
@@
$\left(\mathfrak{B}^{(1)}_{t+1}\right)^{-1}\left[\mathfrak{A}^{(1)}_{t+1}x_{t}+\mathfrak{a}^{(1)}_{t+1}\right$]
@@
$\left(\mathfrak{B}^{(1)}_{t+1}\right)^{-1}\mathfrak{C}_{t+1}\gamma^{*}$
@@
$\mathfrak{Q}_{t+1}=$ $-\left(\mathfrak{A}^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{A}^{(1)}_{t+1}-2\left(\mathfrak{A}^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{C}_{t+1}(\mathfrak{G}_{t+1})^{-1}\mathfrak{A}^{(2)}_{t+1}$ (3.37)
$\mathfrak{q}_{t+1}=$ $-\left(\mathfrak{A}^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{a}^{(1)}_{t+1}-\left(\mathfrak{A}^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{C}_{t+1}(\mathfrak{G}_{t+1})^{-1}\mathfrak{a}^{(2)}_{t+1}$ (3.38)